# On the survival of the long-lived inner disk of PDS 70 Key Words.: accretion, accretion disks – circumstellar matter – planets and satellites: formation – protoplanetary disks1 Introduction

### 11 Mar 2024 v1

@@
PDS 70 remains as the best laboratory to investigate the influence of giant planet formation on the structure of the parental disk. One of the most intriguing discoveries is the detection of a resolved inner disk from ALMA observations, extending up to the orbit of PDS 70b. This inner disk is challenging to explain because most of the dust particles are expected to be trapped at the outer edge of the gap open by PDS 70b and PDS 70c. By performing dust evolution models in combination to radiative transfer simulations that match the gas disk masses obtained from recent thermo-chemical models of PDS 70, we find that when the minimum grain size in the models is larger than 0.1m, there is an efficient filtration of dust particles, and the inner disk is depleted during the first million-year of dust evolution. Therefore, to maintain an inner disk, the minimum grain size in the models needs to be smaller than 0.1m. Only when grains are that small, they are diffused and dragged along with the gas throughout the planets’ gap. The small grains transported in the inner disk grow and drift therein, but the constant reservoir of dust particles that are trapped in the outer edge of the gap and that are continuously fragmenting allows to refill the inner disk on million-year timescales. Our flux predictions at millimetre wavelength of these models agree with ALMA observations. These models predict a spectral index of 3.2 in the outer disk and 3.6 in the inner disk. Our simple analytical calculations show what the inner disk water emission recently observed with JWST may originate from these ice-coated small grains that flow through the gap, grow and drift towards the innermost disk regions, reaching the water snowline. These models may mirror the history and evolution of our Solar System where Jupiter and Saturn played a crucial role shaping the architecture and properties of planets in our Solar System.
On the survival of the long-lived inner disk of PDS 70
@@
Although planets form in protoplanetary disks, their discovery while they are still surrounded by their parental disk remains elusive, and protoplanet candidates are still highly debated (Asensio-Torres et al. 2021; Benisty et al. 2022; Ren et al. 2023). PDS 70 remains as the best laboratory to study the influence of giant planet formation on the disk structure.
PDS 70 is a K7 T Tauri star (5.4 Myr old) located in the Upper Centaurus Lupus association at a distance of 113.47 pc (Müller et al. 2018; Gaia Collaboration et al. 2021). It hosts a disk with a large cavity visible at different wavelengths, including scattered light (Dong et al. 2012) and (sub-) millimeter observations (Hashimoto et al. 2015; Keppler et al. 2019). Inside the cavity, two planets at separations of au (PDS 70b) and au (PDS 70c) have been discovered (Keppler et al. 2018; Müller et al. 2018), which are still accreting material from the disk (with values of for PDS 70b and for PDS 70c, Aoyama & Ikoma 2019; Haffert et al. 2019; Thanathibodee et al. 2019; Hashimoto et al. 2020; Zhou et al. 2021). The masses of these planets are difficult to constrain observationally. Wang et al. (2021) provided dynamical constrains of the planet masses enforcing dynamical stability of the planet orbits, obtaining values of for PDS 70b and for PDS 70c. Most likely, the masses of PDS 70b and PDS 70c range between 1-10 , also supported by hydrodynamical models and radiative transfer simulations of this system (Bae et al. 2019; Toci et al. 2020).
The planetary accretion of PDS 70b and PDS 70c suggests that gas is flowing through the gap, which has been supported by recent observations from The Atacama Large Millimeter/submillimeter Array (ALMA) of different molecules (Keppler et al. 2019; Facchini et al. 2021), where CO and HCO are observed within the orbit of PDS 70b. Portilla-Revelo et al. (2023) analysed these ALMA observations suggesting that the depletion factor of the gas density inside the gap is such that both planets are as massive as . Circumplanetary disks (CPDs) around both planets have been directly or indirectly observed. For PDS 70b, the spectrum in the K-band from the Very Large Telescope/SINFONI supports a scenario with dust around PDS 70b, which was interpreted as a CPD (Christiaens et al. 2019). The CPD of PDS 70c has been directly detected with ALMA (Isella et al. 2019; Benisty et al. 2021). Interesting, tentative co-orbital submillimeter emission within the Lagrangian region L5 of the protoplanet PDS 70b has been recently suggested by re-analysing the ALMA data (Balsalobre-Ruza et al. 2023).
@@
Recently, there have been observational evidence that water vapour is likely delivered to the inner disk because of icy pebbles that are drifting inwards (e.g., Salyk et al. 2011, 2019). Observations of the luminosity of infrared HO emission from Spitzer and the James Webb Space Telescope (JWST) spectra and spatially-resolved dust disk radius obtained from ALMA images, suggest that disks that are large and host sub-structures (that retain the pebbles in the outer disk) have lower HO content (Banzatti et al. 2020, 2023). This is supported by models that include dust evolution, substructures, water sublimation, and the diffusion of water vapour in the inner disk (e.g., Kalyaan et al. 2021, 2023). This is opening a new idea in the field, where the origin of water on the inner terrestrial planets do not only rely on local collisions of water-bearing planetesimals, but where it may mainly come from drifting icy-particles.
JWST observations of PDS 70 with the Mid-InfraRed Instrument (MIRI) revealed a wealth of water lines, implying the presence of water in the inner disk (au) and indicating that potential planets forming therein the inner disk may have access to a water reservoir despite the presence of the two giant planets in the outer disk (Perotti et al. 2023). This challenges current models of dust evolution with giant embedded planets that are massive enough to open deep gaps, such as PDS 70b and PDS 70c, where dust trapping is expected to be very efficient at the edge of their gap.
In this paper, we perform dust evolution and radiative transfer models under the conditions of PDS 70 (Sect. 2) to investigate how its inner disk can survive over million-year time scales. The results of these models and the comparison with current and future ALMA observations are presented in Sect. 3. The results and limitations are discussed in Sect. 4, in addition to the comparison with recent JWST observations. Finally, we present the main conclusions of this work in Sect. 5.
@@ 2.1 Hydrodynamical models
Results from hydrodynamical models of planet-disk interaction performed by Bae et al. (2019) are used as an input for the dust evolution models presented in this work. The models from Bae et al. (2019) are tailored to simulate the structure of PDS 70. They performed these simulations with the two-dimensional, locally isothermal hydrodynamical code FARGO-ADSG (Baruteau et al. 2019), which is an extension of the publicly available FARGO code (Benítez-Llambay & Masset 2016). In the models taken for this work, the mass of PDS 70b is 5 and of PDS 70c is 2.5. The disk viscosity () is assumed constant over radius as well as over time, and equal to . The initial disk mass is .
To input the gas density distribution from the 2D (radial and azimuthal) hydrodynamical simulations to the 1D (radial) dust evolution models, the results from the hydro models are azimuthally averaged over 0.1 Myr (between 0.9 and 1 Myr), after the disk has reached a quasi-steady state (Fig 2 in Bae et al. 2019). After this time, the disk mass is , which is the disk mass that is used as an input in the dust evolution models. The planets migrate in the simulations, so their radial locations change over time, although the migration is pretty slow (0.3 and 0.5 au Myr for b and c, respectively).
The averaged gas surface density and radial grid is used for the dust evolution models, and it is assumed that the gas density remains steady during the dust evolution described in the next subsection.
@@ 2.2 Dust evolution models
For the dust evolution, we use the code Dustpy (Stammler & Birnstiel 2022) version 1.0.3. Dustpy calculates the growth and fragmentation of dust particles, as well as their dynamics. Because we assumed that the gas surface density profile remains constant over time, we do not include gas evolution in these models. Details about the equations solved in these dust evolution models are in Birnstiel et al. (2010) and Stammler & Birnstiel (2022).
@@
The original radial grid from the hydrodynamical simulations that had 672 cells (from 2.2 and 198 au) was interpolated to 100 cells to speed up the dust evolution models. With this resolution, we resolve the outer edge of the gap where particles accumulate with around 26 radial cells, the gap and the inner disk are resolved with higher number of radial cells since the grid is logarithmically spaced. A resolution test is present in Appendix A. The peak of the pressure maximum from the hydro models is around 60 au. This peak should coincide with the peak of the ring observed at millimeter observations at 77 au. In our models we take the pressure maximum as the reference point, and for this reason we scale the results from the hydrodynamical simulation before performing the dust evolution models, such that the pressure maximum is at 77 au. After this correction, the radial grid goes from 2.9 to 249 au.
@@
The stellar parameters are the same as in Bae et al. (2019), this means a 0.85 stellar mass. The viscosity is taken as in the hydrodynamical simulations (), which is also assumed for dust diffusion, vertical and radial turbulence of the dust (Pinilla et al. 2021).
@@
The fragmentation velocity of the particles is assumed to be 10 m s. This fragmentation velocity assumes that the outer layer of the grains is composed by water-ice and for which the sticking force is thought to be strong. Currently, there are no laboratory experiments that reproduce the collisions of water-icy dust particles at the low temperatures and pressures as the ones of protoplanetary disks (Blum et al. 2000; Wada et al. 2009; Gundlach et al. 2018; Arakawa et al. 2023). Laboratory experiments that decreased the temperature of dust collisions experiments down to 150-180 K suggest that water-icy particles may not be as sticky as thought before and that the fragmentation velocity is lower (m s, e.g., Musiolik & Wurm 2019; Steinpilz et al. 2019). Pinilla et al. (2021) demonstrated that if the fragmentation velocity of the particles is m s in the entire disk, the disk needs to have low turbulence to be able to reproduce millimeter observations of typical disks. However, with low turbulence (), a Jupiter mass planet is expected to create multiple type of substructures that are not observed in PDS 70 disk, including multiple rings/gaps and asymmetries (e.g. Ataiee et al. 2013; Bae et al. 2018). For this reason, we decided to work with the hydrodynamical results that assume , for which m s works well to reproduce observations.
@@
The dust density distribution is initially assumed as the interstellar medium to follow a power law as from a minimum grain size, which we vary between [, , ] cm. The maximum grain size in the initial dust size distribution is cm. We will therefore present results from three different models. We note that in Pinilla et al. (2016), we investigated the effect of assuming a different power law for the initial condition of dust distribution on the dust trapping and filtration, and found that the results are not affected by this assumption.
@@
These initial minimum grain sizes are selected to explore the dust filtration at the outer edge of the gap and investigate what size of dust particles are needed to maintain the inner disk. Because the grains are allowed to grow, the maximum grain size of the grid is 10 cm, and we take 7 grid cells for each order of magnitude in mass (Stammler & Birnstiel 2022). This means that the grain size grid has 148, 127, and 113 cells when cm, cm, and cm, respectively. The initial dust-to-gas ratio is 1/100. The volume density of the particles () is set to 0.85g cm, in order to be consistent with the grain composition and opacities assumed in the radiative transfer simulations. The simulations run from 0 to 10 Myr. Under the hypothesis that the two planets required at least 1 Myr to form, in addition to the 1 Myr of planet-disk interaction in the hydrodynamical models, the initial time of the dust evolution models should be seen as at least 2 Myr of the disk age.
@@ 2.3 Radiative transfer simulations
To compare the results from the dust evolution models to previous and future ALMA observations and give predictions for different wavelengths, we perform radiative transfer calculations with RADMC3D (Dullemond et al. 2012). The opacity of each grain size is calculated using optool (Dominik et al. 2021), and assuming that the grains are porous spheres with a vacuum volume fraction of 40%, and composed with 10% silicate, 20% carbon, and 30% water ice (Ricci et al. 2010). This composition gives a volume density of the dust particles of 0.85g cm. Stadler et al. (2022) and Zormpas et al. (2022) demonstrated that with these opacities, there is better agreement between models of dust evolution and radiative transfer with observations, specifically with the millimeter fluxes and spectral indices (in contrast to the DSHARP opacities, Birnstiel et al. 2018).
The only source of radiation is the central star, which is assumed to be a black-body with a temperature of 3972 K. For our calculations, we assume photons and scattering photons.
We calculate the total volume dust density of each grain size as
where and , with being a polar angle. is the dust surface density obtained from the dust evolution models. The radial grid is assumed as in the dust evolution models, and for the vertical and azimuthal grid, we assumed 32 and 64 cells, respectively. The dust scale height for each particle size is given by (Youdin & Lithwick 2007; Birnstiel et al. 2010)
@@ 3.1 Dust density distribution
Figure 1 shows the dust density distribution after 1 Myr (top panes) and 10 Myr (bottom panels) of evolution, as a function of grain size and distance from the star. Note that the y-axis of each column is different because it depends on the initial grain size that is assumed in the models (given in the title of each column). The plots include the St=1 (solid white line), which is proportional to the gas surface density taken from the hydrodynamical simulations.
The left column of Fig. 1 corresponds to models where cm, that is our model with the largest minimum grain size. As expected, in this case, the filtration of dust particles at the outer edge of the gap created by the planets is very efficient. At 1 Myr of evolution, the dust grains that were initially within the gap has drifted towards the star, and after 10 Myr of evolution very little amount of dust is left in the inner disk.
@@
where is the gas radial velocity, the gas pressure, the gas volume density, and the Keplerian frequency.
Only when is positive particles are trapped, this implies that particles bigger than critical size will be filtered out Pinilla et al. (2012), with
Because all the terms in Eq. 5 are quantities related to the gas, except for (which is the volume density of the dust particles set to 1.6 g cm), can be calculated given the gas density distribution from the hydrodynamical simulations. Assuming hydrostatic equilibrium of a locally isothermal disk and viscous accretion to calculate and , respectively; at the outer edge of the gap is m. For this reason, when cm, the dust filtering at the outer edge of the gap is very effective, which leads to a high depletion of dust in the inner disk. In a few words, the dust of the inner disk disconnects from the outer disk, and the dust that was initially inside the gap drifts towards the star, while in the outer disk it drifts towards the pressure maximum at the outer edge of the gap.
@@
The result of efficient filtration is seen in Fig. 2, which corresponds to the total dust disk mass within 16 au as a function of time for the models of Fig. 1. The reference location of 16 au is taken to later compare to the results from Portilla-Revelo et al. (2023). The total dust mass within this inner region at the initial time is the same for all the three models (). In the model where cm, this mass is reduced to (or 0.025 ) after 1 Myr of evolution, and it continues decreasing, reaching values of after 10 Myr.
@@
The results of these models motivate to decrease the size of the initial dust particles as low as the critical grain size (Eq. 5) of 0.1 m or even lower. The middle and right columns of Fig. 1 shows the dust density distribution with cm and cm, respectively. These results show that the dust density distribution in the inner disk increases in these two cases, and that the dust filtration at the outer edge of the gap is not total. The dust mass within 16 au for the cm case decreases to after 1 Myr of dust evolution, similar to the case of cm, but after this time, the dust mass does not significantly decrease over the next Myr (Fig. 2). This is because the inner disk is continuously replenished by the very small grains (0.1m) that can flow through the gap. The replenishment of the inner disk from the outer disk dust trap is more efficient when the initial dust size distribution includes smaller grains, cm, as grains with lower Stokes number can be more efficiently dragged and diffused by the gas. This is reflected in the total amount of dust that remains in the inner disk in these simulations (Fig. 2). For this case, the total amount of dust mass is after 1 Myr of evolution, which remains constant until the end of the simulations (10 Myr). This amount of dust is much lower than the one inferred by Benisty et al. (2021) for the inner disk of 0.08–0.36 , when considering a minimum and a maximum grain size of 0.05 m and 10 m, respectively and the DSHARP opacities (Birnstiel et al. 2018). In our models, the maximum grain size in the inner disk varies between 10-100 m depending on the distance from the star as seen in Fig. 3. Therefore, our models have grains of higher opacity at millimeter wavelengths than those considered in Benisty et al. (2021), especially when comparing the DSHARP opacities with the ones used in our models from Ricci et al. (2010) (see for example Fig. A.3 in Stadler et al. 2022). For this reason, even with lower dust masses, our models can reproduce the millimetre emission of the inner disk observed with ALMA. It is worth to note that the obtained grain distribution in the inner disk (Fig. 3) is very different from the size distribution of interstellar grains from (MRN distribution Mathis et al. 1977) of , which was assumed in the calculations by Benisty et al. (2021), and it is plotted in Fig. 3 for comparison. Most recent dust models for the interstellar medium find a different size distributions than the MRN (e.g. Köhler et al. 2015; Jones et al. 2017; Hensley & Draine 2023), which can also affect the inferred dust mass from the inner disk presented in Benisty et al. (2021).

### 21 Mar 2024 v2

# Survival of the long-lived inner disk of PDS 70 Key Words.: accretion, accretion disks – circumstellar matter – planets and satellites: formation – protoplanetary disks1 Introduction

@@
The K7 T Tauri starPDS 70 remains the best laboratory for investigating the influence of giant planet formation on the structure of the parental disk. One of the most intriguing discoveries is the detection of a resolved inner disk from ALMA observations that extends up to the orbit of PDS 70b. It is challenging to explain this inner disk because most of the dust particles are expected to be trapped at the outer edge of the gap opened by PDS 70b and PDS 70c. By performing dust evolution models in combination with radiative transfer simulations that match the gas disk masses obtained from recent thermo-chemical models of PDS 70, we find that when the minimum grain size in the models is larger than 0.1m, there is an efficient filtration of dust particles, and the inner disk is depleted during the first million year of dust evolution. To maintain an inner disk, the minimum grain size in the models therefore needs to be smaller than 0.1m. Only when grains are that small are they diffused and dragged along with the gas throughout the gap opened by the planets. The small grains transported in the inner disk grow and drift into it, but the constant reservoir of dust particles that are trapped at the outer edge of the gap and that continuously fragment allows the inner disk to refill on million-year timescales. Our flux predictions at millimeter wavelength of these models agree with ALMA observations. These models predict a spectral index of 3.2 in the outer and 3.6 in the inner disk. Our simple analytical calculations show that the water emission in the inner disk that was recently observed with the James Webb Space Telescope may originate from these ice-coated small grains that flow through the gap, grow, and drift toward the innermost disk regions to reach the water snowline. These models may mirror the history and evolution of our Solar System, in which Jupiter and Saturn played a crucial role in shaping the architecture and properties of the planets.
Survival of the long-lived inner disk of PDS 70
@@
Although planets form in protoplanetary disks, it remains hard to discover them while they are still surrounded by their parental disk remains. Protoplanet candidates are still highly debated (Asensio-Torres et al. 2021; Benisty et al. 2022; Ren et al. 2023). The K7 T Tauri star PDS 70 remains the best laboratory for studying the influence of giant planet formation on the disk structure.
PDS 70 is 5.4 Myr old and is located in the Upper Centaurus Lupus association at a distance of 113.47 pc (Müller et al. 2018; Gaia Collaboration et al. 2021). It hosts a disk with a large cavity that is visible at different wavelengths, including scattered light (Dong et al. 2012) and (sub-) millimeter observations (Hashimoto et al. 2015; Keppler et al. 2019). Inside the cavity, two planets at separations of au (PDS 70b) and au (PDS 70c) have been discovered (Keppler et al. 2018; Müller et al. 2018) that are still accreting material from the disk (with values of for PDS 70b and for PDS 70c, Aoyama & Ikoma 2019; Haffert et al. 2019; Thanathibodee et al. 2019; Hashimoto et al. 2020; Zhou et al. 2021). The masses of these planets are difficult to constrain observationally. Wang et al. (2021) provided dynamical constrains of the planet masses by enforcing dynamical stability of the planet orbits. They obtained values of for PDS 70b and for PDS 70c. Most likely, the masses of PDS 70b and PDS 70c range between 1-10 , which is also supported by hydrodynamical models and radiative transfer simulations of this system (Bae et al. 2019; Toci et al. 2020).
The planetary accretion of PDS 70b and PDS 70c suggests that gas flows through the gap. This has been supported by recent observations from The Atacama Large Millimeter/submillimeter Array (ALMA) of different molecules (Keppler et al. 2019; Facchini et al. 2021), where CO and HCO were observed within the orbit of PDS 70b. Portilla-Revelo et al. (2023) analyzed these ALMA observations and suggested that the depletion factor of the gas density inside the gap is such that both planets are as massive as . Circumplanetary disks (CPDs) around both planets have been directly or indirectly observed. For PDS 70b, the spectrum in the K band from the Very Large Telescope/SINFONI supports a scenario of dust around PDS 70b, which was interpreted as a CPD (Christiaens et al. 2019). The CPD of PDS 70c was directly detected with ALMA (Isella et al. 2019; Benisty et al. 2021). Interestingly, tentative co-orbital submillimeter emission within the Lagrangian region L5 of the protoplanet PDS 70b has recently been suggested by reanalyzing the ALMA data (Balsalobre-Ruza et al. 2023).
@@
Observational evidence recently showed that water vapor is likely delivered to the inner disk through icy pebbles that drift inward (e.g., Salyk et al. 2011, 2019). Observations of the luminosity of infrared HO emission from Spitzer and the James Webb Space Telescope (JWST) spectra and the spatially resolved dust disk radius obtained from ALMA images suggest that the disks that are large and host substructures (which retain the pebbles in the outer disk) with a lower HO content (Banzatti et al. 2020, 2023). This is supported by models that include dust evolution, substructures, water sublimation, and the diffusion of water vapor in the inner disk (e.g., Kalyaan et al. 2021, 2023). This introduces a new idea in the field, according to which the origin of water on the inner terrestrial planets does not only rely on local collisions of water-bearing planetesimals, but may mainly come from drifting icy-particles.
The JWST observations of PDS 70 with the Mid-InfraRed Instrument (MIRI) revealed a wealth of water lines, implying the presence of water in the inner disk (au) and indicating that potential planets forming in the inner disk may have access to a water reservoir despite the two giant planets in the outer disk (Perotti et al. 2023). This challenges current models of dust evolution with giant embedded planets that are massive enough to open deep gaps, such as PDS 70b and PDS 70c, where dust trapping is expected to be very efficient at the edge of their gap.
In this paper, we perform dust evolution and radiative transfer models for the conditions of PDS 70 (Sect. 2) to investigate how its inner disk can survive over million-year timescales. The results of these models and the comparison with current and future ALMA observations are presented in Sect. 3. The results and limitations are discussed in Sect. 4, in addition to the comparison with recent JWST observations. Finally, we present the main conclusions of this work in Sect. 5.
@@ 2.1 Hydrodynamical models
The results from hydrodynamical models of planet-disk interaction performed by Bae et al. (2019) were used as input for the dust evolution models presented in this work. The models from Bae et al. (2019) were tailored to simulate the structure of PDS 70. The authors performed these simulations with the two-dimensional locally isothermal hydrodynamical code FARGO-ADSG (Baruteau et al. 2019), which is an extension of the publicly available FARGO code (Benítez-Llambay & Masset 2016). In the models taken for this work, the mass of PDS 70b was 5 and that of PDS 70c was 2.5. The disk viscosity () was assumed to be constant over radius as well as over time, and to be equal to . The initial disk mass was .
To input the gas density distribution from the 2D (radial and azimuthal) hydrodynamical simulations to the 1D (radial) dust evolution models, the results from the hydro models were azimuthally averaged over 0.1 Myr (between 0.9 and 1 Myr) after the disk had reached a quasi-steady state (Fig 2 in Bae et al. 2019). After this time, the disk mass was , which is the disk mass we used as input in the dust evolution models. The planets migrated in the simulations, which means that their radial locations changed over time, although the migration was pretty slow (0.3 and 0.5 au Myr for b and c, respectively).
The averaged gas surface density and radial grid were used for the dust evolution models, and we assumed that the gas density remained steady during the dust evolution described in the next subsection.
@@ 2.2 Dust evolution models
For the dust evolution, we used the code Dustpy (Stammler & Birnstiel 2022), version 1.0.3. Dustpy calculates the growth and fragmentation of dust particles, as well as their dynamics. Because we assumed that the gas surface density profile remained constant over time, we did not include gas evolution in these models. Details about the equations solved in these dust evolution models are given in Birnstiel et al. (2010) and Stammler & Birnstiel (2022).
@@
The original radial grid from the hydrodynamical simulations that had 672 cells (from 2.2 and 198 au) was interpolated to 100 cells to speed up the dust evolution models. With this resolution, we resolved the outer edge of the gap, where particles accumulate with about 26 radial cells, and the gap and inner disk were resolved with more radial cells because the grid is logarithmically spaced. A resolution test is present in Appendix A. The peak of the pressure maximum from the hydro models is about 60 au. This peak should coincide with the peak of the ring observed at millimeter observations at 77 au. In our models, we took the pressure maximum as the reference point, and we therefore scaled the results from the hydrodynamical simulation before performing the dust evolution models, such that the pressure maximum was at 77 au. After this correction, the radial grid extended from 2.9 to 249 au.
@@
The stellar parameters were the same as in Bae et al. (2019), that is, a 0.85 stellar mass. The viscosity was taken as in the hydrodynamical simulations (), which was also assumed for dust diffusion and for the vertical and radial turbulence of the dust (Pinilla et al. 2021).
@@
The fragmentation velocity of the particles was assumed to be 10 m s. This fragmentation velocity assumes that the outer layer of the grains is composed of water-ice, for which the sticking force is thought to be strong. Currently, there are no laboratory experiments that reproduce the collisions of water-ice dust particles at the low temperatures and pressures of those of protoplanetary disks (Blum et al. 2000; Wada et al. 2009; Gundlach et al. 2018; Arakawa et al. 2023). Laboratory experiments that decreased the temperature of dust collision experiments to 150-180 K suggested that water-ice particles may not be as sticky as thought before and that the fragmentation velocity is lower (m s, e.g., Musiolik & Wurm 2019; Steinpilz et al. 2019). Pinilla et al. (2021) demonstrated that when the fragmentation velocity of the particles is m s in the entire disk, the disk needs to have low turbulence to be able to reproduce millimeter observations of typical disks. However, with low turbulence (), a Jupiter-mass planet is expected to create multiple types of substructures, which are not observed in the PDS 70 disk, including multiple rings, gaps, and asymmetries (e.g. Ataiee et al. 2013; Bae et al. 2018). We therefore decided to work with the hydrodynamical results that assumed , for which m s works well to reproduce observations.
@@
The dust density distribution was initially assumed as the interstellar medium to follow a power law as from a minimum grain size, which we varied between [, , ] cm. The maximum grain size in the initial dust size distribution was cm. We therefore present results from three different models. We note that in Pinilla et al. (2016), we investigated the effect of assuming a different power law for the initial condition of dust distribution on the dust trapping and filtration, and we found that the results are not affected by this assumption.
@@
These initial minimum grain sizes were selected to explore the dust filtration at the outer edge of the gap and investigate the size of dust particles that is required to maintain the inner disk. Because the grains were allowed to grow, the maximum grain size of the grid was 10 cm, and we took 7 grid cells for each order of magnitude in mass (Stammler & Birnstiel 2022). This means that the grain size grid had 148, 127, and 113 cells when cm, cm, and cm, respectively. The initial dust-to-gas ratio was 1/100. The volume density of the particles () was set to 0.85g cm in order to be consistent with the grain composition and opacities assumed in the radiative transfer simulations. The simulations ran from 0 to 10 Myr. Under the hypothesis that the two planets required at least 1 Myr to form in addition to the 1 Myr of planet-disk interaction in the hydrodynamical models, the initial time of the dust evolution models should be seen as at least 2 Myr of the disk age.
@@ 2.3 Radiative transfer simulations
To compare the results from the dust evolution models to previous and future ALMA observations and give predictions for different wavelengths, we performed radiative transfer calculations with RADMC3D (Dullemond et al. 2012). The opacity of each grain size was calculated using optool (Dominik et al. 2021), and we assumed that the grains were porous spheres with a vacuum volume fraction of 40% and that they were composed of 10% silicate, 20% carbon, and 30% water ice (Ricci et al. 2010). This composition gives a volume density of the dust particles of 0.85g cm. Stadler et al. (2022) and Zormpas et al. (2022) demonstrated that with these opacities, the models of dust evolution and radiative transfer agree better with observations, specifically, with the millimeter fluxes and spectral indices (in contrast to the DSHARP opacities, Birnstiel et al. 2018).
The only source of radiation is the central star, which was assumed to be a black-body with a temperature of 3972 K. For our calculations, we assumed photons and scattering photons.
We calculated the total volume dust density of each grain size as
where and , with being a polar angle. is the dust surface density obtained from the dust evolution models. The radial grid was assumed as in the dust evolution models, and for the vertical and azimuthal grid, we assumed 32 and 64 cells, respectively. The dust scale height for each particle size was given by (Youdin & Lithwick 2007; Birnstiel et al. 2010)
@@ 3.1 Dust density distribution
Figure 1 shows the dust density distribution after 1 Myr (top panes) and 10 Myr (bottom panels) of evolution as a function of grain size and distance from the star. The y-axis of each column is different because it depends on the initial grain size that is assumed in the models (given in the title of each column). The plots include the St=1 (solid white line), which is proportional to the gas surface density taken from the hydrodynamical simulations.
The left column of Fig. 1 corresponds to models where cm, which is our model with the largest minimum grain size. As expected, the filtration of dust particles at the outer edge of the gap created by the planets is very efficient in this case. At 1 Myr of evolution, the dust grains that were initially within the gap have drifted toward the star, and after 10 Myr of evolution, very little dust is left in the inner disk.
@@
where is the radial gas velocity, is the gas pressure, is the gas volume density, and is the Keplerian frequency.
Only when is positive are particles trapped. This implies that particles larger than critical size are filtered out Pinilla et al. (2012), with
Because all the terms in Eq. 5 are quantities related to the gas, except for (which is the volume density of the dust particles set to 1.6 g cm), can be calculated given the gas density distribution from the hydrodynamical simulations. When we assume hydrostatic equilibrium of a locally isothermal disk and viscous accretion to calculate and , respectively, at the outer edge of the gap is m. For this reason, when cm, the dust filtering at the outer edge of the gap is very effective, which leads to a high depletion of dust in the inner disk. In a few words, the dust of the inner disk disconnects from the outer disk, and the dust that was initially inside the gap drifts toward the star, while in the outer disk, it drifts toward the pressure maximum at the outer edge of the gap.
@@
The result of efficient filtration is seen in Fig. 2, which corresponds to the total dust disk mass within 16 au as a function of time for the models of Fig. 1. The reference location of 16 au is taken below to compare to the results from Portilla-Revelo et al. (2023). The total dust mass within this inner region at the initial time is the same for all the three models (). In the model where cm, this mass is reduced to (or 0.025 ) after 1 Myr of evolution, and it continues to decrease, reaching values of after 10 Myr.
@@
The results of these models motivated us to decrease the size of the initial dust particles to as low as the critical grain size (Eq. 5) of 0.1 m or even lower. The middle and right columns of Fig. 1 show the dust density distribution with cm and cm, respectively. These results show that the dust density distribution in the inner disk increases in these two cases and that the dust filtration at the outer edge of the gap is not total. The dust mass within 16 au for the cm case decreases to after 1 Myr of dust evolution, similar to the case of cm, but after this time, the dust mass does not significantly decrease over the next million years (Fig. 2). This is because the inner disk is continuously replenished by the very small grains (0.1m) that can flow through the gap. The replenishment of the inner disk from the outer disk dust trap is more efficient when the initial dust size distribution includes smaller grains, cm, as grains with lower Stokes number can be more efficiently dragged and diffused by the gas. This is reflected in the total amount of dust that remains in the inner disk in these simulations (Fig. 2). For this case, the total amount of dust mass is after 1 Myr of evolution, which remains constant until the end of the simulations (10 Myr). This amount of dust is much lower than the dust inferred by Benisty et al. (2021) for the inner disk of 0.08–0.36 for a minimum and a maximum grain size of 0.05 m and 10 m, respectively, and the DSHARP opacities (Birnstiel et al. 2018). In our models, the maximum grain size in the inner disk varied between 10-100 m depending on the distance from the star, as shown in Fig. 3. Therefore, our models have grains of higher opacity at millimeter wavelengths than those considered in Benisty et al. (2021), especially when comparing the DSHARP opacities with those used in our models from Ricci et al. (2010) (see for example Fig. A.3 in Stadler et al. 2022). Even with lower dust masses, our models can therefore reproduce the millimeter emission of the inner disk observed with ALMA. It is worthwhile to note that the obtained grain distribution in the inner disk (Fig. 3) is very different from the size distribution of interstellar grains from (MRN distribution Mathis et al. 1977) of , which was assumed in the calculations by Benisty et al. (2021), and this is plotted in Fig. 3 for comparison. Most of the recent dust models for the interstellar medium find a different size distributions than the MRN (e.g. Köhler et al. 2015; Jones et al. 2017; Hensley & Draine 2023), which can also affect the inferred dust mass from the inner disk presented in Benisty et al. (2021).